# A Game-Theoretic Framework for Distributed Load Balancing: Static and Dynamic Game Models

### v1

@@ 6 Conclusions
In this work, we develop a game-theoretic framework for load balancing in static and dynamic settings, motivated by practical applications in job scheduling and cyber-physical systems. We demonstrate that the static game formulation is a potential game, which guarantees the existence of a pure Nash equilibrium, with best-response dynamics converging in a finite number of iterations. To account for the carryover effect, we extend our formulation to dynamic settings and show that players employing best-response strategies in each state can achieve a fully load-balanced system, with convergence to the Nash equilibrium of the static game occurring in polynomial time. Our numerical experiments validate the efficiency of the proposed algorithms.
@@ Appendix A Omitted Proofs
A.1 Proof of Theorem 3.7
Proof A.13.
@@
which is the relative available processing rate of server $j$ that is viewed by player $i$ at time $t=i$. Thus, player $i$ sorts servers in a way that $\hat{\mu}^{t}_{i1}\geq\hat{\mu}^{t}_{i2}\cdots\geq\hat{\mu}^{t}_{im}$. Now suppose player $i$ sends a job to servers indexed in $\{1,\cdots,c_{i}-1\}$. We therefore have for $1\leq j<c_{i}$:
$C^{t}_{i}=\frac{\lambda_{i}a^{t}_{ij}}{\mu_{j}}+\frac{1}{\hat{\mu}^{t}_{ij}},$
where $c_{i}$ is the smallest integer index satisfying
$C_{i}^{t}\leq\frac{1}{\hat{\mu}^{t}_{i,c_{i}}}.$ (25)
Now, it is player $(i+1)$’s turn to update its policy and it calculates the available processing rate for each server. Thus, we have
$\frac{1}{\hat{\mu}^{t+1}_{i+1,j}}=\begin{cases}C_{i}^{t}-\frac{\lambda_{i+1}a^{0}_{i+1,j}}{\mu_{j}}&1\leq j<c_{i},\\[10.0pt] \frac{1}{\hat{\mu}^{t}_{ij}}-\frac{\lambda_{i+1}a^{0}_{i+1,j}}{\mu_{j}}&c_{i}\leq j\leq m.\end{cases}$ (26)
@@
  • $c_{i}-1=m$: This means that player $i$ distributes its job among all servers and we now want to check if player $i+1$ does the same thing. In this case we have
    $C_{i}^{t}=\frac{\lambda_{i+1}+\sum_{j=1}^{m}\frac{\mu_{j}}{\hat{\mu}^{t+1}_{i+1,j}}}{\sum_{j=1}^{m}\mu_{j}}.$
    Suppose server $k$ is the server with highest $\hat{\mu}^{t+1}_{i+1,k}$, so that player $i+1$ sends job to it. Now, we have to check whether player $i$ also sends a job to other servers. Take server $h\in[m]\setminus\{k\}$ as an example. Player $i+1$ sends a portion of its job to server $h$ only if
@@
    which obviously holds because for all $i\in[n],j\in[m]$, $a_{ij}^{0}\neq 0$. More generally, for all $h\in[m]$ and $\mathcal{M}_{i+1}^{h}\subseteq[m]\setminus\{h\}$, we have
@@
    $=$ $\frac{\lambda_{i+1}a^{t-1}_{i+1,h}}{\mu_{h}}+\frac{\lambda_{i+1}}{\sum_{j\in\mathcal{M}_{i+1}^{h}}\mu_{j}}+C_{i}^{t}$ (28)
    -- $\frac{\lambda_{i+1}\sum_{j\in\mathcal{M}_{i+1}^{h}}a^{0}_{i+1,j}}{\sum_{j\in\mathcal{M}_{i+1}^{h}}\mu_{j}}.$ (29)
    Thus, we conclude that player $c_{i+1}-1=m$, or in other words $S^{t}_{i}=S^{t+1}_{i+1}$.
  • $c_{i}-1<m$: Suppose player $i+1$ sends job to servers indexed in $\mathcal{N}^{t+1}_{i+1}\subseteq\{c_{i},\cdots,m\}$. Note that $\mathcal{N}^{t+1}_{i+1}$ can also be an empty set. We now want to see if it sends any job to servers indexed in $S^{t}_{i}$. Take server $k\in S_{i}^{t}$ as an example. Player $i+1$ sends job to server $k$ only if
@@
    Since $a^{0}_{i+1,j}\neq 0$ for all $j\in[m]$, we conclude that
    $\frac{\lambda_{i+1}a^{0}_{i+1,k}}{\mu_{k}}+\frac{\lambda_{i+1}-\lambda_{i+1}\sum_{j\in\mathcal{N}^{t+1}_{i+1}}a^{0}_{i+1,j}}{\sum_{j\in\mathcal{N}^{t+1}_{i+1}}\mu_{j}}>0,$
    and therefore player $i+1$ sends a fraction of its job to server $k$ at time $t+1$. Now, to check whether player $i+1$ also sends a job to the rest of the servers in $S_{i}^{t}$, let us take an arbitrary server $h\in S_{i}^{t}\setminus\{k\}$. Then, player $i+1$ sends a job to server $h$ only if
@@
    $=$ $\frac{\lambda_{i+1}a^{0}_{i+1,h}}{\mu_{h}}$ (33)
    $+$ $\frac{\lambda_{i+1}+C_{i}^{t}\mu_{k}-\lambda_{i+1}a_{i+1,k}^{0}}{\mu_{k}+\sum_{j\in\mathcal{N}^{t+1}_{i+1}}\mu_{j}}$ (34)
@@
    More generally, for any $h\in S_{i}^{t}\setminus\{k\}$ and $\mathcal{H}_{i}^{h}\subseteq S_{i}^{t}\setminus\{h,k\}$, we have
    $C_{i}^{t}<$ $\left(\frac{\sum_{k\in\mathcal{H}_{i}^{h}}\mu_{k}+\sum_{j\in\mathcal{N}^{t+1}_{i+1}}\mu_{j}}{\sum_{j\in\mathcal{N}^{t+1}_{i+1}}\mu_{j}}\right)\left(\frac{\lambda_{i+1}a^{0}_{i+1,h}}{\mu_{h}}\right)$
    $+$ $\lambda_{i+1}\left(\frac{1-\sum_{k\in\mathcal{H}_{i}^{h}}a_{i+1,k}^{0}-\sum_{j\in\mathcal{N}^{t+1}_{i+1}}a^{0}_{i+1,j}}{\sum_{j\in\mathcal{N}^{t+1}_{i+1}}\mu_{j}}\right)$ (38)
@@
    which holds because of (32) and the fact that $a^{0}_{i+1,j}\neq 0$ for all $j\in[m]$. Since $i\in\{1,\cdots,n-1\}$ is chosen arbitrarily, we can conclude that $S^{1}_{1}\subseteq S^{2}_{2}\cdots\subseteq S^{n}_{n}$.
@@
To complete the proof, we note that when player $n$ updates its action, it plays its best response with respect to the action profile $\{a_{1}^{1},a_{2}^{2},\cdots,a_{n-1}^{n-1}\}$. Suppose it sends a portion of its job only to servers indexed in $\{1,\cdots c_{n}-1\}$. We therefore know that $S_{i}^{i}\subseteq\{1,\cdots c_{n}-1\}$ for $i\in\{1,\cdots n-1\}$. Moreover, we know that player $n$ distributes its job in a way such that
$C^{n}_{n}=\frac{\sum_{i=1}^{n}\lambda_{i}a^{i}_{ij}+s_{j}^{0}}{\mu_{j}}\ \ \forall j\in\{1,\cdots m\}.$
In other words, player $n$ ensures that the normalized load is equal across all servers receiving jobs. Consequently, for all players, the servers to which they allocate jobs have identical normalized loads, which is exactly what players achieve when playing their best-response strategies. As a result, no player has an incentive to change their policy, as this distribution aligns with their best-response strategy. Thus, $\{a_{1}^{1},a_{2}^{2},\cdots,a_{n}^{n}\}$ forms a pure NE of the static game.
A.2 An Alternative Bound for the Convergence Rate of the Dynamic Game

### v2

@@ 6 Conclusions
In this work, we develop a game-theoretic framework for load balancing in static and dynamic settings, motivated by practical applications in job scheduling and cyber-physical systems. We demonstrate that the static game formulation is a potential game, which guarantees the existence of a pure Nash equilibrium, with best-response dynamics converging in a finite number of iterations. We then bound the price of anarchy (PoA) of the static game in terms of game parameters. To account for the carryover effect, we extend our formulation to dynamic settings and show that players employing best-response strategies in each state can achieve a fully load-balanced system, with convergence to the Nash equilibrium of the static game occurring in polynomial time. Our numerical experiments validate the efficiency of the proposed algorithms.
@@ Appendix A Omitted Proofs
A.1 Proof of Theorem 7
{pf}
We prove the result by induction on sequential player updates. First, we show that the sets of servers receiving jobs form an increasing sequence as players best respond. Finally, we verify that the last player’s allocation equalizes the normalized loads, ensuring that no player can improve unilaterally.
@@ A.1 Proof of Theorem 7
which is the relative available processing rate of server $j$ that is viewed by player $i$ at time $t=i$. Thus, player $i$ sorts servers in a way that $\hat{\mu}^{t}_{i1}\geq\hat{\mu}^{t}_{i2}\cdots\geq\hat{\mu}^{t}_{im}$. Let $c_{i}$ be the smallest integer index satisfying
$\frac{\lambda_{i}+\sum_{k=1}^{c_{i}-1}\frac{\mu_{k}}{\hat{\mu}^{t}_{ik}}}{\sum_{k=1}^{c_{i}-1}\mu_{k}}\leq\frac{1}{\hat{\mu}^{t}_{i,c_{i}}},$ (26)
and define $C_{i}^{t}\coloneqq\frac{\lambda_{i}+\sum_{k=1}^{c_{i}-1}\frac{\mu_{k}}{\hat{\mu}^{t}_{ik}}}{\sum_{k=1}^{c_{i}-1}\mu_{k}}$. By the best-response rule, player $i$ sends jobs to servers $\{1,\dots,c_{i}-1\}$. For these servers we have
$C_{i}^{t}=\frac{\lambda_{i}a_{ij}^{t}}{\mu_{j}}+\frac{1}{\hat{\mu}_{ij}^{t}}\quad 1\leq j<c_{i}.$
Now when player $i+1$ updates, the inverse available rates satisfy
$\frac{1}{\hat{\mu}^{t+1}_{i+1,j}}=\begin{cases}C_{i}^{t}-\frac{\lambda_{i+1}a^{0}_{i+1,j}}{\mu_{j}}&1\leq j<c_{i},\\ \frac{1}{\hat{\mu}^{t}_{ij}}-\frac{\lambda_{i+1}a^{0}_{i+1,j}}{\mu_{j}}&c_{i}\leq j\leq m.\end{cases}$
@@
  • $c_{i}-1=m$: This means that player $i$ distributes its job across all servers and we now want to check if that is also the case for player $i+1$. Suppose server $k$ is the server with highest $\hat{\mu}^{t+1}_{i+1,k}$, so that we know player $i+1$ sends jobs to it. Now, we have to check whether player $i$ also sends a job to other servers. Take server $h\in[m]\setminus\{k\}$ as an example. Player $i+1$ sends a portion of its job to server $h$ only if
@@
    which holds because we assume that for all $j\in[m]$, $a_{i+1,j}^{0}\neq 0$. More generally, for all $h\in[m]$ and $\mathcal{M}_{i+1}^{h}\subseteq[m]\setminus\{h\}$, we have
@@
    $=$ $\frac{\lambda_{i+1}a^{0}_{i+1,h}}{\mu_{h}}+\frac{\lambda_{i+1}}{\sum_{j\in\mathcal{M}_{i+1}^{h}}\mu_{j}}+C_{i}^{t}$ (28)
    - $\frac{\lambda_{i+1}\sum_{j\in\mathcal{M}_{i+1}^{h}}a^{0}_{i+1,j}}{\sum_{j\in\mathcal{M}_{i+1}^{h}}\mu_{j}},$ (29)
    which holds because for all $i\in[n],j\in[m]$, $a_{i+1,j}^{0}\neq 0$. Thus, we conclude that $c_{i+1}-1=m$, which is equivalent to $S^{t+1}_{i+1}=S^{t}_{i}$.
  • $c_{i}-1<m$: Suppose player $i+1$ assigns jobs to the set of servers $\mathcal{N}^{t+1}_{i+1}\subseteq\{c_{i},\dots,m\}$. Note that $\mathcal{N}^{t+1}_{i+1}$ may also be empty. We now examine whether player $i+1$ assigns any jobs to the servers in $S^{t}_{i}$. Consider, for example, a server $k\in S^{t}_{i}=\{1,\dots,c_{i}-1\}$. Player $i+1$ assigns jobs to server $k$ only if
@@
    Since $a^{0}_{i+1,j}\neq 0$ for all $j\in[m]$, we conclude that player $i+1$ sends a fraction of its job to server $k$ at time $t+1$. To check whether player $i+1$ also assigns jobs to the remaining servers in $S_{i}^{t}$, consider an arbitrary server $h\in S_{i}^{t}\setminus\{k\}$. Player $i+1$ assigns jobs to server $h$ only if
@@
    $=$ $\frac{\lambda_{i+1}a^{0}_{i+1,h}}{\mu_{h}}+\frac{\lambda_{i+1}+C_{i}^{t}\mu_{k}-\lambda_{i+1}a_{i+1,k}^{0}}{\mu_{k}+\sum_{j\in\mathcal{N}^{t+1}_{i+1}}\mu_{j}}$ (33)
@@
    More generally, suppose player $i+1$ sends jobs to servers in $\mathcal{H}_{i+1}^{t+1}\subseteq S_{i}^{t}$. Then for any $h\in S_{i}^{t}\setminus\mathcal{H}_{i+1}^{t+1}$ we have to check
    $C_{i}^{t}<$ $\left(\frac{\sum_{k\in\mathcal{H}_{i+1}^{t+1}}\mu_{k}+\sum_{j\in\mathcal{N}^{t+1}_{i+1}}\mu_{j}}{\sum_{j\in\mathcal{N}^{t+1}_{i+1}}\mu_{j}}\right)\left(\frac{\lambda_{i+1}a^{0}_{i+1,h}}{\mu_{h}}\right)$
    $+$ $\lambda_{i+1}\left(\frac{1-\sum_{k\in\mathcal{H}_{i+1}^{t+1}}a_{i+1,k}^{0}-\sum_{j\in\mathcal{N}^{t+1}_{i+1}}a^{0}_{i+1,j}}{\sum_{j\in\mathcal{N}^{t+1}_{i+1}}\mu_{j}}\right)$ (37)
@@
    which holds because of Eq. (32) and the fact that $a^{0}_{i+1,j}\neq 0$ for all $j\in[m]$. Since $i\in\{1,\cdots,n-1\}$ is chosen arbitrarily, we can conclude that $S^{1}_{1}\subseteq S^{2}_{2}\cdots\subseteq S^{n}_{n}$.
@@
To complete the proof, we note that when player $n$ updates its action, it plays its best response with respect to the action profile $\{a_{1}^{1},a_{2}^{2},\cdots,a_{n-1}^{n-1}\}$. Suppose it sends jobs only to servers indexed in $\{1,\cdots c_{n}-1\}$. We therefore know that $S_{i}^{i}\subseteq\{1,\cdots c_{n}-1\}$ for $i\in\{1,\cdots n\}$. Moreover, we know that player $n$ distributes its job in a way such that
$C^{n}_{n}=\frac{\sum_{i=1}^{n}\lambda_{i}a^{i}_{ij}+s_{j}^{0}}{\mu_{j}}\ \ \forall j\in\{1,\cdots c_{n}-1\}.$ (39)
[…]
A.2 Proof of Theorem 8
In order to prove Theorem 8, we first state and prove the following auxiliary lemma.
Lemma 14.
Given an action profile $a$, we define the normalized load on server $j$ as
$L_{j}(a)\coloneqq\frac{s_{j}^{0}+\sum_{i=1}^{n}\lambda_{i}a_{ij}}{\mu_{j}}.$ (40)
Let $a$ be a pure NE of the static load-balancing game, and denote its support by $S(a)=\{j:a_{ij}>0\text{ for some }i\}$, i.e., the set of servers that receive jobs from at least one player under that NE. Then, the normalized loads on all servers in $S(a)$ must be equal, that is,
$L_{j}(a)=L:=\frac{\sum_{\ell\in S(a)}s_{\ell}^{0}+\sum_{i=1}^{n}\lambda_{i}}{\sum_{\ell\in S(a)}\mu_{\ell}}\quad\forall j\in S(a).$
{pf}
In any pure NE of the static game, each player plays a best response [51], as characterized in Lemma 5. Suppose $a$ is a pure NE. From Eq. (11), for each server $j$ to which player $i$ assigns jobs, we have
$L_{j}(a)$ $=\frac{\lambda_{i}a_{ij}}{\mu_{j}}+\frac{s^{0}_{j}+\sum_{k\neq i}\lambda_{k}a_{kj}}{\mu_{j}}$ (41)
$=\frac{\lambda_{i}+\sum_{k=1}^{c_{i}-1}\tfrac{\mu_{k}}{\hat{\mu}_{ik}}}{\sum_{k=1}^{c_{i}-1}\mu_{k}},$ (42)
where we note that the right-hand side does not depend on $j$. Therefore, in an NE, whenever a player distributes jobs across multiple servers, it equalizes the normalized loads on those servers.
Now suppose in an NE $a$ there exist two players $i$ and $i^{\prime}$, where player $i$ distributes its job only to servers in a set $H$, while player $i^{\prime}$ distributes its job only to servers in a disjoint set $H^{\prime}$. [^6] Otherwise, if $H\cap H^{\prime}\neq\emptyset$, then by the first part of the proof, the normalized loads on all servers in $H\cup H^{\prime}$ must be the same, and there is nothing to prove. From Eq. (41), we know that in any NE, the normalized loads are equal across all servers within the same set. That is,
$\forall j\in H,\;L_{j}(a)=L\quad\text{and}\quad\forall j\in H^{\prime},\;L_{j}(a)=L^{\prime}.$
Our goal is to show that $L=L^{\prime}$. To derive a contradiction, assume without loss of generality that $L^{\prime}<L$. Since player $i$ does not send jobs to any server $j\in H^{\prime}$, using Eq. (12), we obtain:
$\frac{\lambda_{i}+\sum_{k\in H}\frac{\mu_{k}}{\hat{\mu}_{ik}}}{\sum_{k\in H}\mu_{k}}\leq\frac{1}{\hat{\mu}_{ij}}$ $=\frac{s_{j}^{0}+\sum_{h\neq i}\lambda_{h}a_{hj}}{\mu_{j}}$
$=\frac{s_{j}^{0}+\sum_{h}\lambda_{h}a_{hj}}{\mu_{j}}=L^{\prime},$
where the second equality holds because $a_{ij}=0$. On the other hand, we can expand the left-hand side as
$\frac{\lambda_{i}+\sum_{k\in H}\frac{\mu_{k}}{\hat{\mu}_{ik}}}{\sum_{k\in H}\mu_{k}}$ $=\frac{\lambda_{i}+\sum_{k\in H}\left(s_{k}^{0}+\sum_{h\neq i}\lambda_{h}a_{hk}\right)}{\sum_{k\in H}\mu_{k}}$
$=\frac{\sum_{k\in H}\left(s_{k}^{0}+\sum_{h}\lambda_{h}a_{hk}\right)}{\sum_{k\in H}\mu_{k}}$
$=\frac{\sum_{k\in H}\mu_{k}L}{\sum_{k\in H}\mu_{k}}=L,$
where the second equality holds because $\lambda_{i}=\sum_{k\in H}\lambda_{i}a_{ik}$. Thus, we have $L\leq L^{\prime}$, which contradicts the assumption $L^{\prime}<L$. Therefore, we must have $L=L^{\prime}$.
Finally, using the above argument, since the definition of normalized load in (40) is independent of a specific player $i$, we conclude that the normalized loads on all servers that receive jobs from at least one of the players must be equal, that is $L_{j}(a)=L\ \forall j\in S(a)$. Therefore,
$L\sum_{j\in S(a)}\mu_{j}=\sum_{j\in S(a)}\mu_{j}L_{j}(a)=\sum_{j\in S(a)}\left(s_{j}^{0}+\sum_{i=1}^{n}\lambda_{i}a_{ij}\right).$
[…]
$SC(a)$ $=\sum_{i=1}^{n}D_{i}(a)$
$=\sum_{i=1}^{n}\sum_{j=1}^{m}\lambda_{i}a_{ij}\left(\frac{\lambda_{i}a_{ij}}{2\mu_{j}}+\frac{s^{0}_{j}+\sum_{k\neq i}\lambda_{k}a_{kj}}{\mu_{j}}\right).$
Let $a^{\prime}$ be any arbitrary pure Nash equilibrium, and let $a^{*}$ be the action profile that minimizes the social cost. Also define $NE:=SC(a^{\prime})$ and $OPT:=SC(a^{*})$. As in [52], for any $i\in[n]$, we can write
$D_{i}(a^{\prime}_{i},a^{\prime}_{-i})$ $\leq D_{i}(a^{*}_{i},a^{\prime}_{-i})$
$=\sum_{j=1}^{m}\lambda_{i}a^{*}_{ij}\left(\frac{\lambda_{i}a^{*}_{ij}}{2\mu_{j}}+\frac{s^{0}_{j}+\sum_{k\neq i}\lambda_{k}a^{\prime}_{kj}}{\mu_{j}}\right).$
Summing the above inequality for all $i\in[n]$ we get:
$NE$ $=\sum_{i=1}^{n}D_{i}(a^{\prime}_{i},a^{\prime}_{-i})$
$\leq OPT+\sum_{i=1}^{n}\sum_{j=1}^{m}\lambda_{i}a^{*}_{ij}\left(\frac{s_{j}^{0}+\sum_{k=1}^{n}\lambda_{k}a^{\prime}_{kj}}{\mu_{j}}\right).$ (43)
Using the result of Lemma 14, we obtain:
$NE$ $\leq OPT+\sum_{i=1}^{n}\sum_{j=1}^{m}\lambda_{i}a^{*}_{ij}\left(\frac{s_{j}^{0}+\sum_{k=1}^{n}\lambda_{k}a^{\prime}_{kj}}{\mu_{j}}\right)$
$=OPT+L\sum_{i=1}^{n}\sum_{j\in S(a^{\prime})}\lambda_{i}a^{*}_{ij}+\sum_{i=1}^{n}\sum_{j\notin S(a^{\prime})}\lambda_{i}a^{*}_{ij}\left(\frac{s_{j}^{0}}{\mu_{j}}\right)$
$\leq OPT+\max\{L,\max_{j\in[m]}\{\frac{s_{j}^{0}}{\mu_{j}}\}\}\left(\sum_{i=1}^{n}\sum_{j=1}^{m}\lambda_{i}a^{*}_{ij}\right),$
where $L=\frac{\sum_{j\in S(a^{\prime})}s_{j}^{0}+\sum_{i=1}^{n}\lambda_{i}}{\sum_{j\in S(a^{\prime})}\mu_{j}}$. Therefore, we have
$\frac{NE}{OPT}\leq 1+\frac{\max\{L,\max_{j\in[m]}\{\frac{s_{j}^{0}}{\mu_{j}}\}\}\left(\sum_{i=1}^{n}\lambda_{i}\right)}{OPT}.$ (44)
Let us define $X_{j}(a):=\sum_{i=1}^{n}\lambda_{i}a_{ij}$. Then, we can rewrite the social cost of action profile $a$ as
$SC(a)$ $=\sum_{j=1}^{m}\frac{1}{\mu_{j}}\left(\frac{X^{2}_{j}(a)}{2}+s_{j}^{0}X_{j}(a)\right),$
which is a function of $X=(X_{j})_{j\in[m]}$, such that $X\in\mathcal{X}$, where $\mathcal{X}=\{X\in\mathbb{R}_{\geq 0}^{m}:\sum_{j=1}^{m}X_{j}=\sum_{i=1}^{n}\lambda_{i}\}$. Thus,
$OPT$ $=\min_{X\in\mathcal{X}}\sum_{j=1}^{m}\frac{1}{\mu_{j}}\Bigl{(}\frac{X_{j}^{2}}{2}+s_{j}^{0}X_{j}\Bigr{)}\geq\min_{X\in\mathcal{X}}\sum_{j=1}^{m}\frac{X_{j}^{2}}{2\mu_{j}}$
$\geq\min_{X\in\mathcal{X}}\frac{\left(\sum_{j=1}^{m}X_{j}\right)^{2}}{2\sum_{j=1}^{m}\mu_{j}}=\frac{\left(\sum_{i=1}^{n}\lambda_{i}\right)^{2}}{2\sum_{j=1}^{m}\mu_{j}},$
where the last inequality uses the Cauchy–Schwarz inequality. By substituting the above inequality in (44) we get
$\frac{NE}{OPT}\leq 1+2\max\{L,\max_{j\in[m]}\{\frac{s_{j}^{0}}{\mu_{j}}\}\}\left(\frac{\sum_{j=1}^{m}\mu_{j}}{\sum_{i=1}^{n}\lambda_{i}}\right).$ (45)
Finally, we note that
$L=\frac{\sum_{j\in S(a^{\prime})}s_{j}^{0}+\sum_{i=1}^{n}\lambda_{i}}{\sum_{j\in S(a^{\prime})}\mu_{j}}\leq\max_{j\in[m]}\{\frac{s_{j}^{0}}{\mu_{j}}\}+\frac{\sum_{i=1}^{n}\lambda_{i}}{\mu_{\text{min}}}.$ (46)
By substituting (46) into (45), we obtain
$\frac{NE}{OPT}\leq 1+2\left(\max_{j\in[m]}\{\frac{s_{j}^{0}}{\mu_{j}}\}+\frac{\sum_{i=1}^{n}\lambda_{i}}{\mu_{\text{min}}}\right)\left(\frac{\sum_{j=1}^{m}\mu_{j}}{\sum_{i=1}^{n}\lambda_{i}}\right).$
[…]
Corollary A.3.
Suppose $s^{0}_{j}=0\ \forall j$. Then, $PoA\leq 3$.
{pf}
In the special case where $s_{j}^{0}=0$ for all $j$, the support of any Nash equilibrium includes all servers, i.e., $S(a)=m$. In this case, the upper bound in Eq. (46) becomes $L=(\sum_{i=1}^{n}\lambda_{i})/(\sum_{j=1}^{m}\mu_{j})$. Substituting this into Eq. (45), we obtain
$\text{PoA}\leq 1+2\left(\frac{\sum_{i=1}^{n}\lambda_{i}}{\sum_{j=1}^{m}\mu_{j}}\right)\left(\frac{\sum_{j=1}^{m}\mu_{j}}{\sum_{i=1}^{n}\lambda_{i}}\right)=3.$
A.3 An Alternative Bound for the Convergence Rate of the Dynamic Game